# High-order interactions in quantum optomechanics: fluctuations, dynamics and thermodynamics

### v1

@@ I Introduction
The validity of the above-mentioned model relies on two fundamental prerequisites: (i) the oscillation amplitude of any mechanical mode is much smaller than the cavity length, and (ii) the mechanical frequencies are much smaller than all optical frequencies [1]. The first assumption guarantees the absence of nonlinearities caused by broad oscillations of the wall. The second assumption legitimises the rotating wave approximation (RWA), which allows us to safely ignore both secularities and phonon-photon scattering terms in the Hamiltonian [5].
Although these two assumptions are reasonably justified by the state-of-the-art technology of experimental setups [21, 22], recent literature has addressed optomechanical systems in a more extended fashion, namely by including within the Hamiltonian both phonon-photon scattering terms [5, 23, 24, 25, 26] and quadratic terms in the mechanical oscillation amplitudes [1, 27, 17, 28]. Interestingly, this extended formulation of the optomechanical models offers new perspectives and leads to the prediction of new intriguing effects. For example, it has already been shown that the inclusion of scattering terms in the Hamiltonian leads to the prediction phenomena such as vacuum Casimir-Rabi oscillations [29, 30], as well as the conversion of incoherent phonons into correlated photon pairs [31]. Notably, the latter has paved the way for the conception of quantum heat engines based on a Casimir-Otto cycle [32].
@@
The paper is structured as follows: in Section II we review relevant mathematical tools of our analysis, such as the quantization protocol, the generalized master equation, and the thermodynamic quantities of interest. In Section III we introduce the Hamiltonian of the 3-D box, from which we also acquire the Hamiltonian of the 1-D cavity. By means of the perturbation theory, we calculate the correction to the ground state caused by the interaction. Finally, in Section IV we present our numerical analysis, in which we study the time evolution of populations, heat rates and entropy production rates when specific resonance conditions are activated.
II Theoretical background
II.1 Quantizing optomechanical systems
We consider a massless scalar field confined in a cavity as a simple model for the electromagnetic field to obtain a qualitative understanding of the dynamics. The field confinement is imposed by means of boundary conditions to the equations of motion. If one of the cavity walls oscillates, say along the direction $x$, the wall+field system is described by coupled differential equations for both the wall and the field. However, if the oscillation amplitude of the wall $\delta L_{\text{x}}$ is much smaller than the cavity length $L_{\text{x}}$, i.e., $\delta L_{\text{x}}/L_{\text{x}}\ll 1$, it is much simpler to first assume that the wall is fixed, thereby solving the equation of motion for the field with static boundary conditions, and only subsequently introduce a small fluctuation in the wave vector of the field modes.
The procedure to quantize the optomechanical system can be summarized by the following steps:
  1. 1.
    Solve the equations of motion with static boundary conditions for the confined field.
  2. 2.
    Substitute the expressions for the field modes into the Hamiltonian density.
  3. 3.
    Allow for a small fluctuation of the cavity length by replaceing $L_{\text{x}}\rightarrow L_{\text{x}}+\delta L_{\text{x}}$, and then expanding each wave vector $k_{\text{x}}$ in the the Hamiltonian density with respect to the perturbation parameter $\delta L_{\text{x}}/L_{\text{x}}$.
  4. 4.
    Integrate the Hamiltonian density over the box volume.
  5. 5.
    Promote the field mode amplitudes $a_{\mathbb{n}}$ and the small perturbation $\delta L_{\text{x}}$ to the quantum operators $\hat{a}_{\mathbb{n}}$ and $\delta L_{\text{x}}(\hat{b}+\hat{b}^{\dagger})$, respectively.
The last step is complemented by imposing canoncial bosonic commutation rules for the oscillation operators $[\hat{b},\hat{b}^{\dagger}]=1$ as well as for the field modes $[\hat{a}_{\mathbb{n}},\hat{a}_{\mathbb{m}}^{\dagger}]=\delta_{\mathbb{n}\mathbb{m}}$. Note that $\mathbb{n}=(n_{\text{x}},n_{\text{y}},n_{\text{z}})$. We stress that, in this protocol, we expand the Hamiltonian density with respect to $\delta L_{\text{x}}$ only in the wave vectors. This is equivalent to the procedure introduced originally by Law [5], where both the bosonic operators $\hat{a}_{k}$ and the mode frequencies $\omega_{k}$ have been expanded with respect to the oscillation amplitude $x_{m}$ around the average cavity length $l_{0}$.
II.2 Quantum state evolution via master equation
Nonlinear interactions introduce an element of anharmonicity, thereby leading to small modifications in the energy level structure while retaining an overall quasi-harmonic character. As a result, the dynamics of the system has to be described using a generalized master equation formulated without invoking the conventional secular approximation. An appropriate theoretical framework for capturing the time evolution of the density matrix operator $\hat{\rho}$ in hybrid quantum systems subject to dissipation and thermal-like noise has been obtained in previous work [33]. The master equation reads
$\dot{\hat{\rho}}=-i[\hat{H},\hat{\rho}]+\kappa\mathcal{L}[\hat{A}]\hat{\rho}+\gamma\mathcal{L}[\hat{B}]\hat{\rho}\;,$ (1)
where $\kappa$ and $\gamma$ are, respectively, the cavity and mirror damping rates. The Liouvillian superoperator can be expressed in the general form:
$\mathcal{L}[\hat{O}]\hat{\rho}=$ $\frac{1}{2}\sum_{\begin{subarray}{c}j>k\\ l>m\end{subarray}}\{O_{lm}n_{ml}({T})[\hat{P}_{lm}\hat{\rho}\hat{P}_{kj}-\hat{P}_{kj}\hat{P}_{lm}\hat{\rho}]$
$+$ $O_{kj}n_{jk}({T})[\hat{P}_{lm}\hat{\rho}\hat{P}_{kj}-\hat{\rho}\hat{P}_{kj}\hat{P}_{lm}]$
$+$ $O_{lm}[n_{ml}({T})+1][\hat{P}_{kj}\hat{\rho}\hat{P}_{lm}-\hat{\rho}\hat{P}_{lm}\hat{P}_{kj}]$
$+$ $O_{kj}[n_{jk}({T})+1][\hat{P}_{kj}\hat{\rho}\hat{P}_{lm}-\hat{P}_{lm}\hat{P}_{kj}\hat{\rho}]\}\;.$ (2)
The generalized master equation in Eq. (1) (with the dissipators defined as in Eq. (2)) is not of Lindblad form, which implies that essential properties such as the positivity of the density matrix and conservation of probability are not automatically ensured. However, a detailed inspection shows that the dissipator can be regarded as approximately lindblad-like under certain conditions [33]. Specifically, if the energy transition rate $|E_{\beta}-E_{\alpha}|\approx|\omega_{\alpha\beta}|$ is much larger than the system damping rates $\gamma,\kappa$, the corresponding terms contribute negligibly to the dynamics. Therefore, one can assume in Eq. (2) that $n_{jk}({T})\approx n_{ml}({T})$ making the dissipator effectively Lindblad in nature. Despite their limited physical relevance, these quantities result in fast-oscillating terms in the dissipative dynamics that can cause numerical instabilities and significantly increase computational cost.
II.3 Quantum thermodynamics
In the previous section we have presented a method to study the dynamics of the system by means of the generalized master equation. In order to establish an interaction with the environment we assumed that the two subsystems, namely the wall and the cavity field, are coupled to two different baths at different temperatures [53]. This suggests that, if the two subsystems also interact with themselves resonantly, heat will flow from the hot bath to the cold bath, and the whole quantum system is constantly out of equilibrium [54, 55]. To describe the dynamics of the heat flows, we first need to recall some fundamental notion of quantum thermodynamics.
In the absence of external drives, i.e. $\hat{H}istime-independent$, the first law of thermodynamics in its differential form is recovered by computing the time derivative of the average energy $\langle\hat{H}\rangle:=\textrm{Tr}[\hat{H}\hat{\rho}(t)]$ as follows:
where we have introduced the total heat flow $\mathcal{J}(t)=\mathcal{J}_{\text{c}}(t)+\mathcal{J}_{\text{w}}(t)$. Importantly, we define the heat-flow from the cold bath to the cavity by means of the expression $\mathcal{J}_{\text{c}}(t)\equiv\dot{\mathcal{Q}}_{\text{c}}(t)=\textrm{Tr}[\hat{H}\hat{\mathcal{L}}_{\text{c}}\hat{\rho}(t)]$, and the heat-flow from the hot bath to the wall by means of the expression $\mathcal{J}_{\text{w}}(t)\equiv\dot{\mathcal{Q}}_{\text{w}}(t)=\textrm{Tr}[\hat{H}\hat{\mathcal{L}}_{\text{w}}\hat{\rho}(t)]$, see [54, 56].
$\frac{dS}{dt}$ $=-\mathrm{Tr}\left(\frac{d\hat{\rho}}{dt}\ln\hat{\rho}\right)=-\mathrm{Tr}\left[(\mathcal{L}\hat{\rho})\ln\hat{\rho}\right]$
$=-\mathrm{Tr}[\mathcal{L}_{\mathrm{w}}\hat{\rho}\ln\hat{\rho}]-\mathrm{Tr}[\mathcal{L}_{\mathrm{c}}\hat{\rho}\ln\hat{\rho}],$ (6)
$\dot{\Sigma}$ $=-\mathrm{Tr}[\mathcal{L}_{\mathrm{w}}\hat{\rho}\ln\hat{\rho}]-\mathrm{Tr}[\mathcal{L}_{\mathrm{c}}\hat{\rho}\ln\hat{\rho}]$
$\quad-\frac{\mathrm{Tr}(H\mathcal{L}_{\mathrm{w}}\hat{\rho})}{T_{\mathrm{w}}}-\frac{\mathrm{Tr}(H\mathcal{L}_{\mathrm{c}}\hat{\rho})}{T_{\mathrm{c}}}.$ (7)
$-\mathrm{Tr}[\mathcal{L}_{\alpha}\hat{\rho}(\ln\hat{\rho}-\ln\hat{\sigma})]$ $\geq 0,$ (8)
where $\alpha=\mathrm{c}$ or $\mathrm{w}$.
III Highly nonlinear optomechanical model
III.1 Hamiltonian of the 3-D and the 1-D model
@@
To obtain the Hamiltonian of such system we apply the protocol described in Section II.1. However, contrary to what prompted there, we do not stop the perturbative expansion of the Hamiltonian density to first order in $\epsilon$ but we proceed up to the third order. This yields the Hamiltonian
@@
In the formulas above $\Omega$ indicates the oscillation frequency of the wall, whereas the dispersion law of the scalar field reads $\omega_{\mathbb{n}}:=\sqrt{k_{{n_{\textrm{x}}}}^{2}+k_{n_{\perp}}^{2}}$, where we define the wave vectors $k_{{n_{\textrm{x}}}}=\pi n_{\textrm{x}}/L_{\textrm{x}}$, and $k_{n_{\perp}}^{2}=k_{{n_{\textrm{y}}}}^{2}+k_{{n_{\textrm{z}}}}^{2}=\left(\pi n_{\textrm{y}}/L_{\textrm{y}}\right)^{2}+\left(\pi n_{\textrm{z}}/L_{\textrm{z}}\right)^{2}$. Quantum operators fulfill the canonical commutation relations $[\hat{a}_{\mathbb{n}},\hat{a}_{{\mathbb{m}}}^{\dagger}]=\delta_{{\mathbb{n}}{\mathbb{m}}}$ for the field and $[\hat{b},\hat{b}^{\dagger}]=1$ for the harmonic oscillator. We introduced the quadrature position and momentum operators $\hat{X}_{\textrm{b}}=\frac{1}{2}(\hat{b}^{\dagger}+\hat{b})$, $\hat{P}_{\textrm{b}}=\frac{i}{2}(\hat{b}^{\dagger}-\hat{b})$, $\hat{X}_{n}=\frac{1}{2}(\hat{a}_{n}^{\dagger}+\hat{a}_{n})$, and $\hat{P}_{n}=\frac{i}{2}(\hat{a}_{n}^{\dagger}-\hat{a}_{n})$. In our notation $\mathbb{n}\equiv(n_{x},n_{\text{y}},n_{\text{z}})$ and $\mathbb{m}\equiv(m_{x},n_{\text{y}},n_{\text{z}})$ (note that the two sets of indices only differ in the first integer). The explicit expression for $\zeta_{\mathbb{n}\mathbb{m}}$, $\theta_{\mathbb{n}\mathbb{m}}$ and $\chi_{\mathbb{n}\mathbb{m}}$ are reported in Appendix A. We recall that the first order of such Hamiltonian has already been calculated [32], whereas $\hat{H}_{2}$ has recently been already obtained for a 1-dimensional cavity [28].
@@
[…]
@@
[…]
@@
$\hat{H}_{0}:=$ $\sum_{n}\hbar\,\omega_{n}\,\hat{a}_{n}^{\dagger}\hat{a}_{n}+\hbar\,\Omega\,\hat{b}^{\dagger}\hat{b},$
$\hat{H}_{1}:=$ $-4\hbar\sum_{n,m}(-1)^{n+m}\sqrt{\omega_{n}\omega_{m}}\,\hat{X}_{n}\hat{X}_{m}\hat{X}_{b},$
$\hat{H}_{2}:=$ $8\hbar\sum_{n,m}(-1)^{n+m}\sqrt{\omega_{n}\omega_{m}}\,\hat{X}_{n}\hat{X}_{m}\hat{X}_{b}^{2},$
$\hat{H}_{3}:=$ $16\hbar\sum_{n,m}(-1)^{n+m}\sqrt{\omega_{n}\omega_{m}}\left[\frac{k_{n}k_{m}L^{2}}{3}\hat{P}_{n}\hat{P}_{m}\right.$
@@
In the equations above, we defined $L_{\text{x}}\equiv L$ and $\omega_{n}=ck_{n}\equiv n\pi c/L$.
III.2 Perturbed ground state
@@
Figure 1: Eigenenergies of the system as a function of the mechanical frequency (a). The blue lines represent the eigenvalues of the whole Hamiltonian in Eq. (15), whereas the gray lines indicate the eigenvalues calculated from Eq. (15a). Zoom around the first-order resonance $\Omega=2\omega_{1}$ (b), the second-order resonance $2\Omega=2\omega_{1}$ (c), and the third-order resonance $3\Omega=2\omega_{1}$ (d), respectively. Graphs are realized by setting $\epsilon=0.07$.
Figure 2: Time evolution of the populations of the mirror (red line), of cavity mode 1 (blue line), and of cavity mode 2 (green line), fixing the resonant condition $\Omega=2\omega_{1}$. The solid lines represent the dynamics of the populations evolving via the whole Hamiltonian in Eq. (15), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (15a). The two graphs are realized by setting $\epsilon=0.03$ (a), and $\epsilon=0.07$ (b).
Figure 3: Time evolution of $\mathcal{J}_{\text{w}}(t)$ (red line) and $\mathcal{J}_{\text{c}}(t)$ (blue line) after enforcing the resonant condition $\Omega=2\omega_{1}$. The solid lines represent the dynamics of the heat flows evolving via the whole Hamiltonian in Eq. (15), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (15a). The graphs are realized by setting $\epsilon=0.03$ (a), and $\epsilon=0.07$ (b).
Figure 4: Time evolution of the entropy production rate after enforcing the resonant condition. The solid lines represent the dynamics of the populations evolving via the whole Hamiltonian in Eq. (15), whereas the dashed lines indicate the evolution due to the only contribution in Eq. (15a). Plots have been realized by setting $\epsilon=0.03$ (red lines), and $\epsilon=0.07$ (blue lines).
IV Numerical analysis
We are now interested in studying the dynamics and the thermodynamics of our system once specific resonant conditions between the wall and one mode of the confined field are activated. In particular, we investigate the heat transfer between the hot bath coupled to the wall and the cold bath coupled to the cavity by analysing the dynamics of the populations of the involved bosonic degrees of freedom, the rates of heat transferred between modes and baths, as well as the entropy production rate.
Results have been obtained by first numerically solving the master equation in Eq. (1) and subsequently calculating the average values of each quantity of interest. For the realization of all figures below we have used the following parameters: $\omega_{2}=2\omega_{1}=1$, $\gamma=0.018$, $\kappa=0.006$, $T_{\textrm{c}}=10^{-6}$ and $T_{\textrm{w}}=0.3$. Moreover, we have chosen $\kappa_{0}=0.003$ as time scale. Finally, frequencies and temperatures have been normalized with respect to $\omega_{2}$.
IV.1 Three-body interaction
We are interested in considering the resonant interactions between the mechanical mode and two cavity modes $\omega_{1}$ and $\omega_{2}$. It has been shown that the resonant condition with a single cavity mode drastically reduces the effective interactions with the other modes, which can therefore be safely omitted from the dynamics [51]. This reduces the Hamiltonian contributions to $\hat{H}_{0}=\omega_{1}\hat{a}_{1}^{\dagger}\hat{a}_{1}+\omega_{2}\hat{a}_{2}^{\dagger}\hat{a}_{2}+\Omega\hat{b}^{\dagger}\hat{b}$ and
$\hat{H}_{1}:=$ $-4\hbar\,\hat{X}_{\textrm{b}}\left(\omega_{1}\,\hat{X}_{1}^{2}+\omega_{2}\,\hat{X}_{2}^{2}-\sqrt{\omega_{1}\omega_{2}}\,\hat{X}_{1}\hat{X}_{2}\right),$ (15a)
$\hat{H}_{2}:=$ $8\hbar\,\hat{X}_{\textrm{b}}^{2}\left(\omega_{1}\,\hat{X}_{1}^{2}+\omega_{2}\,\hat{X}_{2}^{2}-\sqrt{\omega_{1}\omega_{2}}\,\hat{X}_{1}\hat{X}_{2}\right),$ (15b)
$\hat{H}_{3}:=$ $\frac{(4\pi)^{2}\hbar}{3}\left(\omega_{1}\hat{P}_{1}^{2}+4\omega_{2}\hat{P}_{2}^{2}-2\sqrt{\omega_{1}\omega_{2}}\hat{P}_{1}\hat{P}_{2}\right)\hat{X}_{\textrm{b}}^{3}$
$+\frac{(4\pi)^{2}\hbar}{6}\left(\omega_{1}\hat{X}_{1}^{2}+8\omega_{2}\hat{X}_{2}^{2}-5\sqrt{\omega_{1}\omega_{2}}\hat{X}_{1}\hat{X}_{2}\right)\hat{X}_{\textrm{b}}^{3}$
$-16\hbar\left(\omega_{1}\hat{X}_{1}^{2}-\omega_{2}\hat{X}_{2}^{2}+\sqrt{\omega_{1}\omega_{2}}\hat{X}_{1}\hat{X}_{2}\right)\hat{X}_{\textrm{b}}^{3}$ (15c)

### v2

@@ I Introduction
The validity of the above-mentioned model relies on two fundamental prerequisites: (i) the oscillation amplitude of any mechanical mode is much smaller than the cavity length, and (ii) the mechanical frequencies are much smaller than all optical frequencies [53]. The first assumption guarantees the absence of nonlinearities caused by broad oscillations of the wall. The second assumption legitimises the rotating wave approximation (RWA), which allows us to ignore both secularities and phonon-photon scattering terms in the Hamiltonian [41].
Although these two assumptions are reasonably justified by the state-of-the-art technology of experimental setups [43, 4], recent literature has addressed optomechanical systems in a more extended fashion, namely by including within the Hamiltonian both phonon-photon scattering terms [41, 2, 1, 9, 52] and quadratic terms in the mechanical oscillation amplitudes [53, 42, 7, 11]. Interestingly, this extended formulation of the optomechanical models offers new perspectives and leads to the prediction of new intriguing effects. For example, it has already been shown that the inclusion of scattering terms in the Hamiltonian leads to the prediction phenomena such as vacuum Casimir-Rabi oscillations [17, 46], as well as the conversion of incoherent phonons into correlated photon pairs [67]. Notably, the latter has paved the way for the conception of quantum heat engines based on a Casimir-Otto cycle [20]. It is relevant to notice that the significant advancements in the realization of both high-frequency mechanical oscillators [60, 34, 33, 16] and quantum platforms simulating the dynamical Casimir effect [77, 54, 39, 64] make the observations of these effects promising. We stress that heat transfers mediated by second-order effective wall-wall interactions have already been observed experimentally [23].
@@
The paper is structured as follows: in Section II we introduce the Hamiltonian of the 3-D box, from which we also acquire the Hamiltonian of the 1-D cavity. By means of the perturbation theory, we calculate the correction to the ground state caused by the interaction. Finally, in Section III we present our numerical analysis, in which we study the time evolution of populations, heat rates and entropy production rates when specific resonance conditions are activated.
II Highly nonlinear optomechanical model
II.1 Hamiltonian of the 3-D and the 1-D model
@@
To obtain the Hamiltonian of such system we apply the protocol described in [21]. Such protocol is reported in Appendix A.1. This yields the Hamiltonian
@@
[…]
We recall that the first order of such Hamiltonian has already been predicted in [20], however, we want to spend a few words on the role of the first term in $\hat{H}_{1}$. First, we notice that such term does not contribute to the phonon-photon conversion, as the presence of wave vector components $k_{n_{\perp}}$ would not fulfill the momentum conservation in the direction y and z (phonons are created only along x), but it only affects the so-called radiation pressure term. This reads
$\hat{H}_{\textrm{RP}}=2\sum_{\mathbb{n}}\frac{k_{n_{\perp}}^{2}-k_{n_{x}}^{2}}{\omega_{\mathbb{n}}}\left(\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{n}}+\frac{1}{2}\right)\hat{X}_{\textrm{b}}.$ (3)
To provide a physical origin of this term, we note that inside the cavity each photon must have components in each direction due to the Heisenberg principle: the bounded space of the box increases the probability to find the photon in the box volume, thereby also increasing the uncertainty of the photon momentum. On the contrary, outside the box the field modes are not spatially bound and virtual photons pressing on the wall can safely possess only the component along x, as the infinite space outside can ensure high moment precision. Therefore, the presence of terms proportional to $k_{n_{\perp}}$ stems from the difference between the force pushing the wall from the inside the box, caused by photons whose momentum is subject to high uncertainty, as well as the force caused by photons pressing from outside.
@@
  • Resonant interactions. In the interaction Hamiltonians $\hat{H}_{2}$ and $\hat{H}_{3}$ one finds resonant interactions, which consist of four-wave-mixing-like phenomena described by terms of the form $\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{m}}^{\dagger}\hat{b}^{2}+\textrm{h.c.}$ or $\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{m}}\hat{b}^{2}+\textrm{h.c.}$ and $\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{m}}^{\dagger}\hat{b}^{3}+\textrm{h.c.}$ or $\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{m}}\hat{b}^{3}+\textrm{h.c.}$ respectively. Such terms describe scattering processes between two photons and two or three phonons, and they play an active role in the dynamics only if the frequency of the wall fulfills resonant conditions of the form $2\Omega=\omega_{\mathbb{n}}+\omega_{\mathbb{m}}$, $2\Omega=\omega_{\mathbb{n}}-\omega_{\mathbb{m}}$, $3\Omega=\omega_{\mathbb{n}}+\omega_{\mathbb{m}}$ or $3\Omega=\omega_{\mathbb{n}}-\omega_{\mathbb{m}}$.
  • Counterrotating terms. The contributions called counterrottaing terms are $\hat{a}_{\mathbb{n}}\hat{a}_{\mathbb{m}}\hat{b}^{2}+\textrm{h.c.}$ and $\hat{a}_{\mathbb{n}}\hat{a}_{\mathbb{m}}\hat{b}^{3}+\textrm{h.c.}$ As already noted for the counterrotating terms in $\hat{H}_{1}$ [41, 21], the contribution of such terms concerns only high-order interactions. In the current scenario, counterrotating terms play a role in processes where the effective Hamiltonian is at least proportional to $\sim\epsilon^{3}$.
@@
  • Kerr-like interactions. The Hamiltonian $\hat{H}_{3}$ describes the interaction of an odd number of bosons in a similar fashion to $\hat{H}_{1}$. For this reason, we cannot expect static terms (that commute with the free Hamiltonian) in the interaction picture. Nevertheless, we can isolate contributions that display Kerr-like interactions where the interaction itself is mediated by the presence of phonons or photons. Examples of off-resonant terms of this type are $\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{n}}\hat{b}^{\dagger}\hat{b}(\hat{b}+\hat{b}^{\dagger})$, where the radiation pressure is determined by both the number of photons and phonons. Furthermore, one can also find resonant terms, i.e., those where resonant conditions can lead to time independent coefficients, and they read $\hat{b}^{\dagger}\hat{b}\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{m}}^{\dagger}\hat{b}+\textrm{h.c.}$ or $\hat{b}^{\dagger}\hat{b}\hat{a}_{\mathbb{n}}^{\dagger}\hat{a}_{\mathbb{m}}\hat{b}+\textrm{h.c.}$. Note that resonance conditions required to “activate” these terms are the same as those predicted for the interaction Hamiltonian $\hat{H}_{1}$.
@@
$\hat{H}_{0}:=$ $\sum_{n}\,\omega_{n}\,\hat{a}_{n}^{\dagger}\hat{a}_{n}+\Omega\,\hat{b}^{\dagger}\hat{b}$,
$\hat{H}_{1}:=$ $-4\sum_{n,m}(-1)^{n+m}\sqrt{\omega_{n}\omega_{m}}\,\hat{X}_{n}\hat{X}_{m}\hat{X}_{b},$
$\hat{H}_{2}:=$ $8\sum_{n,m}(-1)^{n+m}\sqrt{\omega_{n}\omega_{m}}\,\hat{X}_{n}\hat{X}_{m}\hat{X}_{b}^{2},$
$\hat{H}_{3}:=$ $16\sum_{n,m}(-1)^{n+m}\sqrt{\omega_{n}\omega_{m}}\left[\frac{k_{n}k_{m}L^{2}}{3}\hat{P}_{n}\hat{P}_{m}\right$.
@@
In the equations above, we defined $L_{\text{x}}\equiv L$ and $\omega_{n}=ck_{n}\equiv n\pi c/L$. We note that $\hat{H}_{2}$ has recently been obtained for a 1-dimensional cavity [11].
II.2 Perturbed ground state
@@
Figure 1: Eigenenergies of the system as a function of the mechanical frequency (a). The blue lines represent the eigenvalues of the whole Hamiltonian in Eq. (8), whereas the gray lines indicate the eigenvalues calculated from Eq. (8b). Zoom around the first-order resonance $\Omega=2\omega_{1}$ (b), the second-order resonance $2\Omega=2\omega_{1}$ (c), and the third-order resonance $3\Omega=2\omega_{1}$ (d), respectively. Graphs are realized by setting $\epsilon=0.07$.
III Numerical analysis
The main goal of this section is to examine the effects of high-order interactions encoded in the interaction Hamiltonians $\hat{H}_{2}$ and $\hat{H}_{3}$ on both the dynamics and the thermodynamics of the system. To do so, henceforth we restrict our analysis to the three-body interactions between the wall and two cavity modes, $\omega_{1}$ and $\omega_{2}$, focusing on the resonant interactions between the mechanical mode and the cavity mode $\omega_{1}$. Although in a previous work we have already demonstrated that the presence of the off-resonant mode $\omega_{2}$ do not contribute significantly to the dynamics [19], we cannot exclude a priori that the presence of the high-order terms $\hat{H}_{2}$ and $\hat{H}_{3}$ may alter our previous results.
III.1 Three-body interaction Hamiltonian
The Hamiltonian of the optomechanical system describing the interaction of two cavity modes and the quantum mechanical degree of freedom of the wall is $\hat{H}=\hat{H}_{0}+\epsilon\hat{H}_{1}+\epsilon^{2}\hat{H}_{2}+\epsilon^{3}\hat{H}_{3}$, where
$\hat{H}_{0}:=$ $\,\omega_{1}\hat{a}_{1}^{\dagger}\hat{a}_{1}+\omega_{2}\,\hat{a}_{2}^{\dagger}\hat{a}_{2}+\Omega\,\hat{b}^{\dagger}\hat{b}$ (8a)
$\hat{H}_{1}:=$ $-4\,\hat{X}_{\textrm{b}}\left(\omega_{1}\,\hat{X}_{1}^{2}+\omega_{2}\,\hat{X}_{2}^{2}-\sqrt{\omega_{1}\omega_{2}}\,\hat{X}_{1}\hat{X}_{2}\right),$ (8b)
$\hat{H}_{2}:=$ $8\,\hat{X}_{\textrm{b}}^{2}\left(\omega_{1}\,\hat{X}_{1}^{2}+\omega_{2}\,\hat{X}_{2}^{2}-\sqrt{\omega_{1}\omega_{2}}\,\hat{X}_{1}\hat{X}_{2}\right),$ (8c)
$\hat{H}_{3}:=$ $\frac{(4\pi)^{2}}{3}\left(\omega_{1}\hat{P}_{1}^{2}+4\omega_{2}\hat{P}_{2}^{2}-2\sqrt{\omega_{1}\omega_{2}}\hat{P}_{1}\hat{P}_{2}\right)\hat{X}_{\textrm{b}}^{3}$
$+\frac{(4\pi)^{2}}{6}\left(\omega_{1}\hat{X}_{1}^{2}+8\omega_{2}\hat{X}_{2}^{2}-5\sqrt{\omega_{1}\omega_{2}}\hat{X}_{1}\hat{X}_{2}\right)\hat{X}_{\textrm{b}}^{3}$
$-16\left(\omega_{1}\hat{X}_{1}^{2}-\omega_{2}\hat{X}_{2}^{2}+\sqrt{\omega_{1}\omega_{2}}\hat{X}_{1}\hat{X}_{2}\right)\hat{X}_{\textrm{b}}^{3}$ (8d)